OpenStudy (anonymous):

Choose the correct simplification of (2xy^2)^2(y^2)^3.

OpenStudy (mayliz):

(2xy^2)^2(y^2)^3=4x^2y^10

OpenStudy (anonymous):

((2*x*y^2)^2)*((y^2)^3)\[\left(4 x^2 y^4\right)*\left(y^6\right) \]\[4 x^2 y^{10} \]